Shift report 20010920 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Sep 20 2001 - 08:28:09 EDT

  • Next message: Betty McBreen: "[Fwd: DiskSpace]"

    ------------------------------------------------------------
    Date: Thu Sep 20 08:28:08 2001, Shift: 00:00-08:00
    Supervisor: J.I.Jordre, with: NN
    Log book#9 pages 13 - 14
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam most of the shift. Only about 1 hour interruption to
       fill RHIC from 03:53 to 04:55. However, still only about 50%
       of the ZDC rates of the other experiments. Called MCR twice
       asking for steering, but no increased rates.
    2. Run 4817 (started last shift) with 346k trigger 1 events
       Run 4818                      with  63k trigger 1 events
       Run 4822                      with 254k trigger 1 events
       Run 4823 (still running)      with  80k trigger 1 events
       => Now beyond 1M trigger 1 events at current magnet settings.
       Should therefore change FS to 4kG B polarity at the end of this
       store.
    3. YK tinkered with the ZDC low gain ADC Blue Right module 1 cable
       in the FEH, i.e. unplugging and replugging the cable. That
       solved the problem reported in last evening shift.
       T5 anode 1 and all anodes in T3 tripped once, without particularly
       high currents. Voltages brought back up to nominal values without
       any problem.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Sep 20 2001 - 08:29:19 EDT